7.Moon dust may sometimes be levitated by the charging of dust particles by various means. What electric field is required to levitate dust with a charge of +1e, radius of 20 microns, and weight of 8.2x10^-11 N?
Answer #1

F = qE = m g

mg = W = 8.2 * 10^-11 N

qE = 1 * 1.602 * 10^-19

1.602 * 10^-19 * E = 8.2 * 10^-11

E = 5.1186 *10^8 N/C